OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)]

OpenStudy (bahrom7893):

g(-2) = -12-1=-13 f(-13) = -2*-13-1=26-1=25

OpenStudy (precal):